OpenStudy (anonymous):

How many different arrangements of seven letters can be made using the letters in the name "ULYSSES"?

Parth (parthkohli):

Don't answer him please. He's one of the fiends.

OpenStudy (anonymous):

fiends?

OpenStudy (anonymous):

What do you mean?

Parth (parthkohli):

Answer fiends are the one who need hw help, and just are concerned with the answer.

OpenStudy (anonymous):

Oops... I gave him 2 answers :(

OpenStudy (anonymous):

lol

Parth (parthkohli):

Hmm. If someone doesn't respond and closes the question without even thanking, then don't answer him/her.

Parth (parthkohli):

My work is done here. Thank you.

OpenStudy (anonymous):

we not in it for thanks or medals :P

OpenStudy (anonymous):

if you could tell all 7 letters apart it would be \(7!\) but since there are 3 indistinguishable "s" it is \(\frac{7!}{3!}\)

Parth (parthkohli):

Please don't go further, satellite.

OpenStudy (anonymous):

compute via \[\frac{7!}{3!}=7\times 6\times 5\times 4\] there are a lot of rules around here these days

OpenStudy (anonymous):

Thank you so much and I am not a fiend whoever said that I am failing math right now and I need to do well on this take home test so please I just need answers.

OpenStudy (anonymous):

yw

Parth (parthkohli):

You need to work it out. Please. This community is not the one to ask for answers.

Parth (parthkohli):

It's our duty to explain it to you.

OpenStudy (anonymous):

ul lu 2 --- uly uyl yul ylu luy lyu 6 ---- i think the pattern will be n! just ULYSSES 7! right?

Parth (parthkohli):

7! over 6 @timo86m

OpenStudy (anonymous):

oh the indistinguishable s :)